OpenStudy (idealist10):

Solve y'=(2x+y+1)/(x+2y-4) implicitly.

OpenStudy (idealist10):

@myininaya

OpenStudy (idealist10):

So dy/dx=(2x+y+1)/(x+2y-4), and then?

OpenStudy (idealist10):

Please help me.

ganeshie8 (ganeshie8):

is it homogeneous ?

ganeshie8 (ganeshie8):

if not, can you make it homegeneous forcibly ?

ganeshie8 (ganeshie8):

because you can always solve a homogeneous equation

OpenStudy (idealist10):

y=ux u=y/x y'=u'x+u y'=(2+y/x+1/x)/(1+2(y/x)-4/x) =(1/x)(2x+y+1)/(1/x)(x+2y-4) =(2x+y+1)/(x+2y-4)

ganeshie8 (ganeshie8):

making a linear equation homogeneous is easy, all you need to is to shift it up/down : 

ganeshie8 (ganeshie8):

that substitution works only with homogeneous equations @Idealist10 so think of a way to change the given equation into homogeneous form

ganeshie8 (ganeshie8):

those constants +1 and -4 are making this equation difficult, right ?

OpenStudy (idealist10):

Yes.

ganeshie8 (ganeshie8):

lets get rid off them y'=(2x+y+1)/(x+2y-4) say \(\large u = x+h\) \(\large v = y+k\) \[\large \dfrac{dy}{dx} = \dfrac{dv}{du} \] the given equation changes to : \[\large \dfrac{dv}{du} = \dfrac{2(u-h) + v-k+1}{u-h+2(v-k)-4} \]

ganeshie8 (ganeshie8):

choose the values of "h" and "k" such that they eat off the constant terms

OpenStudy (inkyvoyd):

a system of equations?

ganeshie8 (ganeshie8):

simplifying a bit you get \[\large \dfrac{dv}{du} = \dfrac{2u + v + (\color{red}{-2h-k+1})}{u+2v + (\color{red}{-h-2k-4})} \]

ganeshie8 (ganeshie8):

exactly setting those constant terms equal to 0 gives a system of two linear equations which we can solve for good h,k values

ganeshie8 (ganeshie8):

it seems \(\large h=2, k=-3 \) are the magic numbers that choice of h,k kills the constant terms and you're left with your favorite homogeneous equation to solve : \[\large \dfrac{dv}{du} = \dfrac{2u + v}{u+2v } \]

ganeshie8 (ganeshie8):

solve it, we can change back to xy coordinates after solving this

OpenStudy (idealist10):

@myininaya